The rawfonts package

Alan Jeffrey

v0.01

Overview

The rawfonts L^{A}TEX 2ε package is used to provide emulation of L^{A}TEX 2.09 documents which used low-level commands such as \tenrm. If you say:

    \usepackage{rawfonts}

then over sixty fonts such as \tenrm will be loaded into L^{A}TEX. This has a large overhead, for example this document uses:

    8221 words of font info for 30 fonts

but when run with the rawfonts package it uses:

    19294 words of font info for 73 fonts

That is rawfonts can cause your document to load twice as many fonts. (This overhead is why L^{A}TEX 2ε does not define \tenrm and friends by default.)

If you want to only load a small number of fonts, you can use the only option, for example to only load \tenrm and \tensf:

    \usepackage[only,tenrm,tensf]{rawfonts}

The rawfonts package is intended mainly for use with L^{A}TEX 2.09 documents, and you might want to have the package loaded every time you use L^{A}TEX 2ε in compatibility mode. In this case, you should add the line:

    \RequirePackage{rawfonts}

to your latex209.cfg L^{A}TEX 2.09 compatibility configuration file.